### v1

# Streaming Federated Learning with Markovian Data

Abstract
Federated learning (FL) is now recognized as a key framework for communication-efficient collaborative learning. Most theoretical and empirical studies, however, rely on the assumption that clients have access to pre-collected data sets, with limited investigation into scenarios where clients continuously collect data. In many real-world applications, particularly when data is generated by physical or biological processes, client data streams are often modeled by non-stationary Markov processes. Unlike standard i.i.d. sampling, the performance of FL with Markovian data streams remains poorly understood due to the statistical dependencies between client samples over time. In this paper, we investigate whether FL can still support collaborative learning with Markovian data streams. Specifically, we analyze the performance of Minibatch SGD, Local SGD, and a variant of Local SGD with momentum. We answer affirmatively under standard assumptions and smooth non-convex client objectives: the sample complexity is proportional to the inverse of the number of clients with a communication complexity comparable to the i.i.d. scenario. However, the sample complexity for Markovian data streams remains higher than for i.i.d. sampling.
1 Introduction
As edge networks grow in scale, clients are increasingly diverse [Lim et al., 2020]. Many of these clients are capable of continuously collecting and learning from data, instead of simply storing large quantities of historical data [Khodadadian et al., 2022]. This streaming data arises in applications ranging from health [Greco et al., 2019, Guo et al., 2024] and environmental monitoring [Hill and Minsker, 2010], to control of robots [Kober et al., 2013].
Each data stream provides detailed information about individual clients. However, applications such as environmental resource management [Sirimewan et al., 2024] or public health monitoring [Chae et al., 2018] often aim to derive population-level insights, leveraging data streams from multiple individuals to learn a global model. As this data may be sensitive or costly to communicate, centralized training solutions are undesirable. Federated Learning (FL) has emerged as a key strategy for communication-efficient collaborative learning [Kairouz et al., 2019]. In conventional FL scenarios, clients have access to local datasets used to train models locally. These models are then transmitted to a central server, where aggregation facilitates collaborative learning. FL minimizes communication overhead while learning models with high accuracy and ensuring client data are not shared.
Clients with streaming data must also cope with memory constraints [Marfoq et al., 2023, Liu et al., 2023]. This means that clients have limited control over data sampling: local updates can only be computed using the current data samples stored in memory. Moreover, data in health and environmental monitoring applications is often generated by non-linear physical and biological systems. The data generated by these systems is frequently modeled as non-stationary Markovian data streams [Anderson and Kurtz, 2015, Van Kampen, 1992]. When client memory caches are continually refreshed by Markovian streaming data, the samples available to compute updates are statistically dependent and non-stationary in time.
In this paper, we address the following question: can FL with Markovian data streams support collaborative learning? While Markovian sampling has been explored in the centralized setting [Even, 2023], existing work in the federated setting has focused on Federated Reinforcement Learning [Qi et al., 2021b, Khodadadian et al., 2022, Mangold et al., 2024] with convex objectives. However, classification and regression tasks in systems with non-linear dynamics often rely on deep neural networks [Yuan et al., 2020, Lucia et al., 2020], resulting in non-convex minimization problems during training. We therefore consider a family of training tasks beyond the work in [Qi et al., 2021b, Khodadadian et al., 2022, Mangold et al., 2024] characterized by $L$-smooth non-convex loss functions. Our main contributions can be summarized as follows:
  1. (i)
    Impact of Markovian sampling: We rigorously characterize the impact of Markovian sampling on the convergence rate of Minibatch and Local stochastic gradient descent (SGD), the two most common baselines in FL. A key conclusion is that linear speed-up is achieved under standard assumptions on the objectives and heterogeneity: the sample complexity is proportional to the inverse of the number of clients.
  2. (ii)
    Mitigating client heterogeneity: We show that a momentum-based variant of Local SGD introduced in the i.i.d. setting [Cheng et al., 2024, Yan et al., 2025] can effectively mitigate the client drift with Markovian data streams. In particular, unlike local SGD, local SGD with momentum matches the performance of Minibatch SGD up to a constant without requiring any bound on client data heterogeneity.
2 Related Works
Edge devices often encounter the challenge of processing continuous data streams while operating under memory constraints [Marfoq et al., 2023, Liu et al., 2023]. Learning from streaming data presents numerous challenges with solutions dependent on the data-generating process. One line of research addresses scenarios where the data distribution evolves over time, with continual learning methods designed to adapt the model to the changing data distribution [Schwarz et al., 2018]. Another line focuses on adversarial settings, where the data stream is generated by an adversary [Shalev-Shwartz, 2012], often addressed using online learning algorithms with regret guarantees.
In this paper, we focus on learning from data generated by an unknown Markovian process. Markovian processes have been successfully applied to model various physical and biological systems [Anderson and Kurtz, 2015, Van Kampen, 1992]. Markovian data streams encompass the special case where data samples are drawn i.i.d. from an unknown distribution.
In the centralized setting, this learning problem can be framed as a stochastic approximation problem, typically addressed by variants of stochastic gradient descent (SGD). While SGD with i.i.d. data is well-understood [Robbins and Monro, 1951, Lan, 2020], the Markovian nature of the data stream introduces challenges due to the temporal correlation between samples.
For uniformly ergodic Markov chains, Duchi et al. [2011] first demonstrated that mirror descent achieves optimal convergence rates for Lipschitz, general convex, and non-convex problems. Recently, Dorfman and Levy [2022] proposed a random batch size algorithm that adapts to the mixing time of the Markov chain, while Beznosikov et al. [2023] adopted the same technique to develop accelerated methods and established lower bounds for strongly convex objectives.
For general Markov chains, existing analysis of SGD-type algorithms [Sun et al., 2018, Doan et al., 2020b, a] have shown sub-optimal dependence on the mixing time and variance of stochastic gradient estimates relative to the lower bounds in Beznosikov et al. [2023], Duchi et al. [2011]. We refer to Beznosikov et al. [2023], Table 1 for an exhaustive review of the literature on SGD with Markovian noise.
In the case of SGD with Markovian noise on a finite state space—a common scenario in distributed optimization [Bjorn Johansson, 2007]Even [2023], Hendrikx [2023] developed convergence theories and variance reduction techniques for first-order methods, demonstrating improvements over the gossip algorithm [Boyd et al., 2006].
In the context of FL, a few papers have considered learning from data streams. Odeyomi and Zaruba [2021] and Yoon et al. [2021] framed the problem in an adversarial and in a continual learning setting, respectively. Chen et al. [2020] imposed strong homogeneity assumptions, such as requiring all clients to share a common optimal model. Marfoq et al. [2023] assumed data heterogeneity across clients but i.i.d. data streams at each client and considered the effect of heterogeous memory constraints. Liu et al. [2023] studied how to jointly tune batch sizes and number of local updates, but do not have convergence results in the streaming setting.
In the case of Markovian data streams, recent work has primarily focused on Federated Reinforcement Learning (FRL) with linear function approximation. A key challenge is demonstrating that collaboration is beneficial in this setting; namely, that the per-client sample complexity of FRL algorithms decreases inversely with the number of clients. While this linear speed-up is well-understood for i.i.d. data, the impact of collaboration remains an open question in the presence of statistical dependence, as is the case with Markovian data streams. Khodadadian et al. [2022], Fabbro et al. [2023] demonstrated the benefits of cooperation under a strong homogeneity assumption on the client environment. Heterogeneous settings have been considered in Doan et al. [2019], Jin et al. [2022] with a linear speed-up established in Wang et al. [2024] subject to heterogeneity assumptions, which were then relaxed in Mangold et al. [2024] by incorporating control variates [Karimireddy et al., 2020].
Beyond Markovian data streams, temporal correlations can arise from other factors in the FL setting. For instance, memory update strategies, as explored in Marfoq et al. [2023], can introduce dependencies between subsequent updates. Similarly, the availability of clients can exhibit Markovian behavior, as discussed in Ribero et al. [2023], Rodio et al. [2024], Sun et al. [2025]. While these sources of correlation are distinct from the Markovian data stream focus of this paper, they highlight the broader significance of understanding and mitigating temporal dependencies in the FL context.
3 Problem setup
3.1 Streaming Federated Learning
Consider $M$ clients, each with an objective given by
$F_{m}(w):=\mathbb{E}_{x\sim\pi_{m}}[f_{m}(w;x)],$ (1)
where $f_{m}:\mathbb{R}^{d}\times\Omega_{m}\rightarrow\mathbb{R}$ is a smooth function on $\mathbb{R}^{d}$ and $\pi_{m}$ is the target data distribution for client $m$. In the context of supervised learning, the objective $F_{m}$ corresponds to the true risk of the loss function $f_{m}$, parameterized by the parameter $w$, on a data sample $x$ drawn from $\pi_{m}$.
In FL, the $M$ clients collaborate to solve
$\min_{w\in\mathbb{R}^{d}}\frac{1}{M}\sum_{m=1}^{M}F_{m}(w)$ (2)
by iterating a two-phase procedure over multiple communication rounds. Two popular examples of this procedure are Minibatch SGD (Algorithm 2 in the Appendix) and Local SGD or FedAvg (Algorithm 3 in the Appendix) [Kairouz et al., 2019]. In the first phase of the $t$-th communication round, client $m$ has access to $K$ data samples $(x_{t}^{(m,k)})_{k\in[K]}$, which are used to compute a local update, corresponding to the gradient estimate $g_{t}^{m}$ in Minibatch SGD or iterate $w_{t}^{(m,K)}$ in Local SGD. In the second phase, the updates are aggregated via averaging.
In the standard FL scenario, the samples $(x_{t}^{(m,k)})_{k\in[K]}$ available at client $m$ in any communication round $t$ are fixed, corresponding to pre-collected data [Kairouz et al., 2019]. In contrast, in streaming FL, the data $(x_{t}^{(m,k)})_{k\in[K]}$ available to client $m$ can change in every communication round. Streaming FL models scenarios where clients have a limited memory cache of $K$ samples and new data is collected over time [Marfoq et al., 2023, Liu et al., 2023].
Streaming FL raises a key challenge: the data sampling process is not completely controlled by the clients. As a consequence, i.i.d. data samples from the target distribution $\pi_{m}$ may not be available. For example, patient data in health monitoring is continuously collected from wearable devices. Statistical dependence between data collected over time naturally arises due to latent factors such as daily routines, sleep patterns, and weather conditions [Li et al., 2017]. Moreover, the initial data collection or warm-up period [Baghdadi et al., 2021] might capture atypical behavior, such as adjustments to new wearable devices. As a consequence, the data stream of a patient $m$ will in general be non-stationary and not reflect the long-term target distribution $\pi_{m}$.
3.2 Markovian Data Streams
Data streams arising in health and environmental monitoring often exhibit specific statistical structures. Indeed, many biological, chemical, and physical systems are governed by non-stationary Markov processes [Anderson and Kurtz, 2015, Van Kampen, 1992]. Data streams arising in federated reinforcement learning also form Markov processes [Khodadadian et al., 2022, Qi et al., 2021a]; however, the use of stationary exploration policies leads to stationary data streams. stationary Markovian data stream.
Let $x_{t}^{m}=(x_{t}^{(m,k)})_{k\in[K]}$ be the data samples available to client $m$ in communication round $t$. We model the data stream $X_{m}=(x_{t}^{m},\leavevmode\nobreak\ t\in\mathbb{N})$ of client $m$ by a time-homogeneous Markov chain evolving on the state space $\Omega_{m}\subseteq\mathbb{R}^{d}$ with the corresponding Borel $\sigma$-field $\mathcal{B}_{m}$ Roberts and Rosenthal [2004], Meyn et al. [2009]. The evolution of the time-homogeneous Markov chain for client $m$ is characterized by the initial distribution $x_{0}^{m}\sim Q_{m}$ and by the transition kernel $P_{m}(x_{t}^{(m,k)},\mathrm{d}x_{t}^{(m,k+1)})$. The probability of transitioning from the state $x_{t}^{(m,k)}\in\Omega_{m}$ to a state $x_{t}^{(m,k+1)}\in\mathcal{X}_{m}\in\mathcal{B}_{m}$ is given by $\int_{x\in\mathcal{X}_{m}}P_{m}(x_{t}^{(m,k)},\mathrm{d}x)$. The $t$-step transition kernel for client $m$ from an initial state $x_{0}^{m}$ is denoted by $P_{m}^{t}(x_{0}^{m},\mathrm{d}x)$.
We focus on the case of independent clients, where $X_{m}$ is independent of $X_{m^{\prime}}$ for $m\neq m^{\prime}$. This scenario occurs when client data streams are generated by non-interacting processes. Key examples include patient health monitoring and environmental monitoring across spatially separated regions, where observations from different clients are often assumed to be independent [Sainani, 2010]. [^1] In the case of nearby or overlapping regions, statistical dependence between the data samples may arise Montero et al. [2015].
The Markov chain $X_{m}$ admits a stationary distribution $\pi_{m}$ on $\Omega_{m}$ if
$\int_{x\in\Omega_{m}}\pi_{m}(\mathrm{d}x)P_{m}(x,\mathrm{d}x_{t}^{(m,k)})=\pi_{m}(\mathrm{d}x_{t}^{(m,k)}).$ (3)
We say that $X_{m}$ is stationary if $x_{0}^{m}\sim\pi_{m}$; otherwise, $X_{m}$ is non-stationary.
The stationary distributions $\pi_{m},\leavevmode\nobreak\ m\in[M]$ correspond to the target distributions in (2), which capture the long-term statistics of the data samples. Since long-term statistics, rather than transient behavior of the samples, are of interest for learning in health and environmental monitoring scenarios, the stationary distribution is a natural choice for the target distribution. In order to guarantee that the Markov data stream associated with each client admits a stationary distribution, we impose the following assumption.
Assumption 3.1.
The data samples of client $m\in[M]$ are drawn from an independent time-homogeneous Markov chain $X_{m}$ defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$ with transition kernel $P_{m}$ and initial distribution $\mathcal{Q}_{m}$, converging to the unique stationary distribution $\pi_{m}$.
Assumption 3.1 ensures that samples $x_{t}^{m}$ from the data stream of client $m$ will be approximately drawn from $\pi_{m}$ as $t$ diverges. As each client data stream evolves independently, the system-level data stream $X=((x_{t}^{m})_{m\in[M]},\leavevmode\nobreak\ t\in\mathbb{N})$ will eventually draw samples from the stationary distribution $\pi$ defined on $(\Omega:=\bigtimes_{m=1}^{M}\Omega_{m},\mathcal{B}:=\bigotimes_{m=1}^{M}\mathcal{B}_{m})$.
Let $Q:=\bigotimes_{m=1}^{M}Q_{m}$, $P(x_{0},\mathrm{d}x):=\prod_{m=1}^{M}P_{m}(x_{0}^{m},\mathrm{d}x^{m})$, and denote the $t$-step transition kernel for the system-level Markov chain initially in state $x_{0}$ by $P^{t}(x_{0},\mathrm{d}x)$. In the case $Q\neq\pi$, the system-level data stream is non-stationary. The distance of the distribution of $(x_{t}^{m})_{m\in[M]}$ and $\pi$ can be quantified by the mixing time of the system-level data stream $X$ in terms of the total variation norm.
Definition 3.2.
For two probability measures $\mu,\nu$ defined on $\left(\Omega,\mathcal{B}\right)$, the total variation distance between $\mu$ and $\nu$ is defined as:
$\left\lVert\mu-\nu\right\rVert_{TV}:=2\underset{A\in\mathcal{B}}{\sup}\left\lvert\mu(A)-\nu(A)\right\rvert.$ (4)
Definition 3.3.
For the Markov chain defined on $(\Omega,\mathcal{B})$ with transition kernel $P$ and unique stationary distribution $\pi$, let $d(t):=\underset{x\in\Omega}{\sup}\|P^{t}(x,\cdot)-\pi(\cdot)\|_{\mathrm{TV}}$. Then, the mixing time of the chain is defined as:
$\tau(\epsilon)$ $:=\inf\{t\geq 1:d(t)\leq\epsilon\}.$ (5)
We denote $\tau(1/4)$ simply as $\tau$ and, following convention, we call it the mixing time of the Markov chain.
Following prior work on stochastic optimization with Markovian data [Beznosikov et al., 2023, Nagaraj et al., 2020, Even, 2023], we assume that the system-level Markov chain is uniformly ergodic. This assumption ensures that the system-level Markov data stream admits exponential mixing, i.e., regardless of the initial distribution, it converges exponentially fast to the stationary distribution $\pi$.
Assumption 3.4.
The system-level Markov chain $X$ has a finite mixing time $\tau$. Moreover, it is uniformly ergodic, that is, for every initial distribution $Q$ and $t\in\mathbb{N}$:
$\begin{split}\underset{(x_{1},\dots,x_{M})\in\Omega}{\sup}&\left\lVert QP^{t}((x_{1},\dots,x_{M}),.)-\pi(.)\right\rVert_{\mathrm{TV}}\\ &\leq\left(\frac{1}{4}\right)^{\lfloor t/\tau\rfloor}.\end{split}$ (6)
4 Convergence analysis
4.1 Assumptions
To conduct the convergence analysis, we first impose the following assumptions on the objective functions and the stochastic gradients, which allow for non-convexity.
Assumption 4.1.
The global objective function $F$ is $L$-smooth; that is, for all $w_{1},w_{2}\in\mathbb{R}^{d}$:
$F(w_{2})$ $\leq F(w_{1})+\langle\nabla F(w_{1}),w_{2}-w_{1}\rangle$ (7)
$+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (8)
In some cases, we will need the following assumption that requires the sample-wise local objective functions to be $L$-smooth:
Assumption 4.2.
For every $m\in[M]$, the sample-wise objective functions $f_{m}(w;x)$ are $L$-smooth; that is, for all $w_{1},w_{2}\in\mathbb{R}^{d},x\in\Omega_{m}$,
$f_{m}(w_{2};x)$ $\leq f_{m}(w_{1};x)+\langle\nabla f_{m}(w_{1};x),w_{2}-w_{1}\rangle$ (9)
$+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (10)
We note that Assumption 4.2 implies Assumption 4.1.
Next, we state an assumption on the noise of the stochastic gradients.
Assumption 4.3.
For all clients $m\in[M]$, for all $x\in\Omega_{m}$, and $w\in\mathbb{R}^{d}$, there exists $\sigma>0$ such that
$\left\lVert\nabla f_{m}(w;x)-\nabla F_{m}(w)\right\rVert^{2}\leq\sigma^{2}.$ (11)
Assumption 4.3 uniformly bounds the gradient estimation error for each client. While stronger than the bounded variance assumption, this assumption is standard for stochastic optimization with Markovian noise in both the centralized [Beznosikov et al., 2023] and federated [Mangold et al., 2024] settings.
Heterogeneity in client objective functions and data distributions is a well-known difficulty in FL algorithms based on local SGD, which leads to client drift [Karimireddy et al., 2020]. In order to provide convergence guarantees in the i.i.d. setting, it is necessary to impose constraints on the local gradient norms [Woodworth et al., 2020, Yun et al., 2022]. Client heterogeneity poses the same difficulty in the Markovian setting. As such, we impose the following bounded gradient dissimilarity (BGD) assumption.
Assumption 4.4.
There exist $\theta\geq 0$ and $\delta\geq 1$ such that, for all $w\in\mathbb{R}^{d}$,
$\frac{1}{M}\sum_{m=1}^{M}\left\lVert\nabla F_{m}(w)\right\rVert^{2}\leq\theta^{2}+\delta^{2}\left\lVert\nabla F(w)\right\rVert^{2}.$ (12)
This assumption was first introduced in [Karimireddy et al., 2020] and has since become ubiquitous in the analysis of local SGD [Yun et al., 2022]. In Section 4, we establish iteration and communication complexity bounds of local SGD under these assumptions.
Finally, we define the following three classes of problems for the analysis of minibatch SGD, local SGD, and local SGD-M. Recall from Sec. 3.2 that the system-level Markov chain is denoted by $X$. We assume throughout that all objective functions are in general non-convex and $F$ is bounded from below by $F^{*}>-\infty$.
[…]
We note that
$\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)\subset\mathcal{F}_{2}(L,\sigma,\tau)\subset\mathcal{F}_{1}(L,\sigma,\tau).$ (19)
In the following sections, we analyze the convergence of Minibatch SGD, Local SGD, and Local SGD-M for the problem classes $\mathcal{F}_{1},\mathcal{F}_{3}$, and $\mathcal{F}_{2}$, respectively. We establish conditions, for any $\epsilon>0$, on the step sizes, the number of local steps $K$, and the number of communication rounds $T$ in order to guarantee
$\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2},$ (20)
where $\hat{w}_{T}$ is the output of each algorithm, drawn uniformly at random from the iterates $w_{0},\dots,w_{T-1}$.
We use the Landau big-O notation $\mathcal{O}(\cdot)$, $a\vee b$ for $\max\left\{a,b\right\}$, $a\wedge b$ for $\min\left\{a,b\right\}$, $\Delta_{0}$ for $F(w_{0})-F^{*}$ and $G_{0}$ for $\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left[\left\lVert\nabla F_{m}(w_{0})\right\rVert^{2}\right]$.
4.2 Minibatch SGD
We first establish the following upper bound for Minibatch SGD.
Theorem 4.5.
For the problem class $\mathcal{F}_{1}(L,\sigma,\tau)$, with global step size $\gamma\leq 1/L$, the iterates of Minibatch SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\gamma T}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (21)
The proof of Theorem 4.5 is given in Appendix D. The first deterministic term depends on the initial iterate and decays with the same rate as in the i.i.d setting. The last two terms arise from the dependence structure and non-stationarity of the Markov data process. In contrast to i.i.d. sampling [Woodworth et al., 2020], these terms cannot be controlled by the step size. In particular, the second term reveals that the stochastic gradient noise is amplified by the mixing time of the system-level Markov chain, and the last term is due to the bias of Markovian sampling whenever the system-level Markov chain is non-stationary.
The following corollary characterizes the communication and sample complexity of Minibatch SGD.
Corollary 4.6.
Under the conditions of Theorem 4.5, the required number of local steps and communication rounds for Minibatch SGD to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are:
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right).$ (22)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (23)
The proof of Corollary 4.6 follows directly from Theorem 4.5 and is given in Appendix D.
4.3 Local SGD
In the analysis of Local SGD, bounded heterogeneity assumptions are required, even in the i.i.d. setting [Karimireddy et al., 2020]. We consider the problem class $\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, which requires sample-wise smoothness (Assumption 4.2) and bounded heterogeneity (Assumption 4.4).
Theorem 4.7.
For any $F\in\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, with local step size $\eta\leq\mathcal{O}\left(\frac{1}{LK\delta^{2}}\right)$, the iterates of Local SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\eta KT}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{LK\eta\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (24)
The proof of Theorem 4.7 is given in Appendix E. The convergence behavior of Local SGD closely resembles that of Minibatch SGD, except for the additional third term introduced by the client drift. We note that this term can be controlled by having a sufficiently small step size. Similar to Corollary 4.6, the following corollary specifies the communication and sample complexity of Local SGD for both stationary and non-stationary Markov chains.
Corollary 4.8.
Under the conditions of Theorem 4.7 and
$\eta\leq\mathcal{O}\left(\frac{\delta^{2}\epsilon^{2}}{KL(\theta^{2}+\sigma^{2})}\wedge\frac{1}{KL\delta^{2}}\right),$ (25)
the required number of local steps and communication rounds for Local SGD to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$, are:
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (26)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (27)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (28)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (29)
The proof of Corollary 4.8 is given in Appendix E.
4.4 Local SGD with Momentum
In the case of i.i.d sampling, several algorithms have been proposed to mitigate the client drift effect without the need of Assumption 4.4. One line of work focuses on the control variate technique, first introduced by Karimireddy et al. [2020]. Recently, the work by Cheng et al. [2024], Yan et al. [2025] has demonstrated that momentum can also be used to mitigate client drift in the i.i.d. setting.
In this section, we study the convergence of a momentum-based variant of Local SGD, named Local SGD-M, detailed in Algorithm 1. The key difference from Local SGD is that the local updates are computed via the convex combination of the gradient estimate for the local objective and the aggregated updates from the previous communication round. We show that, with Markovian data, this technique also mitigates the impact of client drift. In particular, the lower bound on communication complexity in the i.i.d. setting [Patel et al., 2022] is achieved up to a constant without any heterogeneity assumptions.
Algorithm 1 Local SGD-M
  Input: initial model $w_{0}$ and gradient estimate $v_{0}$, local learning rate $\eta$, global learning rate $\gamma$ and momentum $\beta$
  for $t=0$ to $T-1$ do
     for every client $m\in[M]$ in parallel do
        Initialize local model $w_{t}^{(m,0)}=w_{t}$
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           $v_{t}^{m,k}=\beta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)+\left(1-\beta\right)v_{t}$
           $w_{t}^{(m,k+1)}=w_{t}^{(m,k)}-\eta v_{t}^{(m,k)}$
        end for
        Communicate $w_{t}^{m,K}$
     end for
     Aggregate: $v_{t+1}=\frac{1}{\eta MK}\sum_{m=1}^{M}\left(w_{t}-w_{t}^{(m,K)}\right)$
     Server update: $w_{t+1}=w_{t}-\gamma v_{t+1}$
  end for
  Output: $\hat{w}_{T}$ sampled uniformly from $w_{0},\dots,w_{T-1}$.
The convergence result for Local SGD with momentum holds for the broader function class $\mathcal{F}_{2}(L,\sigma,\tau)$, which retains the sample-wise smoothness condition but relaxes the requirement for bounded heterogeneity.
Theorem 4.9.
For the problem class $\mathcal{F}_{2}(L,\sigma,\tau)$, with the following conditions on the step sizes:
$\begin{split}\eta KL\leq\mathcal{O}&\left(\frac{1}{\beta}\wedge\frac{1}{\sqrt{MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\right.\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \left.\wedge\frac{1}{\beta\gamma LT}\right),\\ \gamma\leq\mathcal{O}&\left(\frac{\beta}{L}\right),\end{split}$ (30)
the iterates of Local SGD-M satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (31)
The proof of Theorem 4.9 is given in Appendix F.6. The communication and sample complexities of Local SGD-M are characterized in the following corollary.
Corollary 4.10.
Under the conditions of Theorem 4.9, the required number of local steps and communication rounds for Local SGD-M to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are:
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right).$ (32)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (33)
The proof of Corollary 4.10 is given in Appendix F.7.
4.5 Discussion
Table 1 summarizes the communication and sample complexities of Minibatch and Local SGD, and the momentum-based variant of Local SGD, for both stationary and non-stationary Markovian data streams.
The communication complexity is measured by the total number of communication rounds, while the sample complexity refers to the total number of samples processed. For our FL algorithms, the sample complexity is equal to $KT$ and can be estimated by multiplying the bounds for $K$ and $T$ given in Corollaries 4.6, 4.8, and 4.10.
In the stationary case, at the cost of more local computation, Minibatch SGD can achieve the lower bound in Patel et al. [2022], while Local SGD suffers from client drift, and can only reach this lower bound in the low-heterogeneity low-noise regime; i.e., when $\frac{\theta^{2}+\sigma^{2}}{\epsilon^{2}}\leq\delta^{4}$.
In the case of a non-stationary system-level Markov chain, an additional term $\frac{\sigma^{2}}{\epsilon^{2}}$ is introduced in the communication complexity of both Minibatch and Local SGD. Convergence can only be established in the high-precision regime (small $\epsilon$). Minibatch SGD can only achieve the lower bound in Patel et al. [2022] if the gradient noise is sufficiently small $\sigma^{2}\leq\mathcal{O}(L\Delta_{0})$.
Local SGD-M successfully mitigates the client drift caused by heterogeneity. In particular, the lower bound on the communication complexity in Patel et al. [2022] is achieved without any heterogeneity assumptions, up to a constant of $1/\beta$, while maintaining the same number of samples per communication round $K$. This contrasts with SCAFFLSA [Mangold et al., 2024], which utilizes control variates to mitigate heterogeneity, resulting in lower communication complexity. However, SCAFFLSA needs a higher number of samples per communication round than its vanilla version FedLSA and still relies on a heterogeneity assumption around the optimum.
[…]
Table 1: Comparison of Algorithms in Terms of Communication and Sample Complexity
Communication $T$ Sample complexity $KT$
Lower bound i.i.d. [Patel et al., 2022] $\Omega\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\Omega\left(\frac{L\Delta_{0}\sigma}{M\epsilon^{3}}\right)$
MAG [Dorfman and Levy, 2022]* $\tilde{\mathcal{O}}\left(\tau_{m}\frac{(M+L+G)^{2}G^{2}}{\epsilon^{4}}\right)$
Stationary Minibatch SGD $\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{M\epsilon^{4}}\right)$
Local SGD $\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{M\epsilon^{4}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$
Local SGD-M $\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{\beta M\epsilon^{4}}\right)$
Non-stationary Minibatch SGD $\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{\left(L\Delta_{0}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\right)$
Local SGD $\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ $\mathcal{O}\left(\tau\frac{\left(L\Delta_{0}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$
Local SGD-M $\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{\left(\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\right)$
    Here $\tau_{m}$ is the mixing time of client $m$’s Markov chain, $M$ and $G$ are upper bounds on the domain diameter and on the gradient norm, respectively, and $\tilde{\mathcal{O}}$ ignores logarithmic constants.
5 Numerical Results
(a) $K=10$
(b) $K=100$
(c) $K=1000$
Figure 1: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD, and Local SGD-M, with 10 clients, $\gamma=0.01,\eta=0.001$, $\beta=0.1$, and the mixing time for every client $\tau_{m}=100$.
(a) Local SGD
(b) Minibatch SGD
(c) Local SGD-M
Figure 2: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, $\tau_{m}=100$ for all $m$ and $K=10$.
In this section, we experimentally evaluate the performance of Minibatch SGD, local SGD, and local SGD-M on a non-convex linear regression problem with Markovian data. In the following experiments, we study the average performance over $10$ random seeds in a setting similar to Dorfman and Levy [2022].
Let $\mathcal{U}\left(a,b\right)$ denote the uniform probability distribution over $[a,b]$. The data stream for each client $m$ is generated by a two-state Markov chain $(i_{t}^{(m,k)},\leavevmode\nobreak\ k\in[K],t\in\mathbb{N})$, where the probability of jumping from one state to another is $p\in\left(0,1\right)$, with mixing time $\Theta(1/p)$. [^2] Here we use the standard Big-$\Theta$ notation. Associated with each state $i\in\{0,1\}$ are the vectors $V_{m,i}\in\mathbb{R}^{10}$ which are drawn randomly for each seed according to $\mathcal{U}\left(0,1\right)$. For each seed and $i\in\{0,1\}$, half of the optimal parameters $w_{m,i}\in\mathbb{R}^{10}$ take the value $w_{i}^{1}\sim\mathcal{U}(0,1)$, while the other half take the value $w_{i}^{2}\sim\mathcal{U}(1,2)$.
The samples associated with client $m$ correspond to the observations
$x_{t}^{(m,k)}=w_{m,i_{t}^{(m,k)}}^{T}V_{m,i_{t}^{(m,k)}}+\epsilon_{t}^{(m,k)},$ (34)
where $\epsilon_{t}^{(m,k)}\sim\mathcal{U}(0,0.01),\leavevmode\nobreak\ t\in\mathbb{N},m\in[M],k\in[K]$. Since the Markov chain $(i_{t}^{(m,k)},\leavevmode\nobreak\ k\in[K],t\in\mathbb{N})$ is symmetric, the stationary distribution for every client $m$ is uniform; i.e., $\pi_{m}(0)=\pi_{m}(1)=1/2$. The local objective function for each client $m$ is then given by
$F_{m}(w)=$ $\frac{1}{2}\left[\left(w-w_{m,1}\right)^{T}V_{m,1}\right]^{2}$ (35)
$+\frac{1}{2}\left[\left(w-w_{m,2}\right)^{T}V_{m,2}\right]^{2}+\lambda r(w),$
where $r(w)=\frac{1}{2}\sum_{d=1}^{10}\frac{w_{[d]}^{2}}{1+w_{[d]}^{2}}$ is a non-convex regularizer often used in robust non-convex and smooth stochastic optimization [Tran et al., 2021, Fatkhullin et al., 2023].
Impact of Heterogeneity: In Figure 1, we plot the trajectories of the gradient norm $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ over the communications rounds for each of the three algorithms, with different number of samples per communication round. As predicted by our analysis, the performance of Local SGD stalls when the number of local updates increases, while Minibatch SGD and Local SGD-M are more robust to the heterogeneity with improved performance as the number of samples $K$ in each communication round is increased.
Impact of Number of Clients $M$: In Figure 2, we plot the trajectories of the gradient norm $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for each of the three algorithms, with different numbers of clients. The gradient norm for all three algorithms decreases when the number of clients increases, indicating that collaboration is beneficial even with heterogeneous Markovian data.
6 Conclusion
Statistical dependence is a reality in data streams arising from physical and biological systems. A key question is whether collaboration remains beneficial for FL with Markovian data streams. To address this question, we showed that there is a speed-up in FL with smooth non-convex objectives for Minibatch SGD, Local SGD, and Local SGD with Momentum. However, there is a cost associated with Markovian data streams, quantified by an increase in the sample complexity compared with i.i.d. sampling.
Appendix A Preliminaries
A.1 Pseudo-code for Minibatch SGD and Local SGD
Algorithm 2 Minibatch SGD
  Input: initial model $w_{0}$, global learning rate $\gamma$.
  for $t=0$ to $T-1$ do
     for every client $m\in[M]$ in parallel do
        Receive the global model $w_{t}$
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           Compute $g_{t}^{(m,k)}=\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)$
        end for
        Communicate $g_{t}^{m}=\frac{1}{K}g_{t}^{m,k}$
     end for
     Aggregate: $g_{t}=\frac{1}{M}\sum_{m=1}^{M}g_{t}^{m}$
     Server update: $w_{t+1}=w_{t}-\gamma g_{t}$
  end for
  Output: $\hat{w}_{T}$ sampled uniformly from $w_{0},\dots,w_{T-1}$.
Algorithm 3 Local SGD
  Input: initial model $w_{0}$, local learning rate $\eta$.
  for $t=0$ to $T-1$ do
     for every client $m\in[M]$ in parallel do
        Initialize the local model $w_{t}^{(m,0)}=w_{t}$
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           $w_{t}^{(m,k+1)}=w_{t}^{(m,k)}-\eta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)$
        end for
        Communicate $w_{t}^{(m,K)}$
     end for
     Server update: $w_{t+1}=\frac{1}{M}\sum_{m=1}^{M}w_{t}^{(m,K)}$
  end for
  Output: $\hat{w}_{T}$ sampled uniformly from $w_{0},\dots,w_{T-1}$.
A.2 Useful notation
The following notation will be utilized throughout the appendix.
$\bar{g}_{t}=\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right),$ (36)
$g_{t}=\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right),$ (37)
$\Delta_{t}=\mathbb{E}\left[F\left(w_{t}\right)\right]-F^{*},$ (38)
$s_{t}=\mathbb{E}\left[\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\right].$ (39)
We use $\sum_{m,k}$ for $\sum_{m=1}^{M}\sum_{k=1}^{K}$ unless explicitly stated differently.
With this notation, the update rule of Minibatch SGD is
$w_{t+1}=w_{t}-\gamma\bar{g}_{t},$ (40)
the update rule of Local SGD is
$w_{t+1}=w_{t}-\tilde{\eta}_{t}g_{t},$ (41)
where $\tilde{\eta}_{t}=K\eta_{t}$, and the update rule of Local SGD-M is
$w_{t+1}=w_{t}-\gamma v_{t+1}.$ (42)
Appendix B Markov chain Preliminaries
We recall the definition of total variation distance:
Definition B.1.
For two probabilities $\mu,\nu$ defined on $\left(\Omega,\mathcal{B}\right)$, the total variation distance between $\mu$ and $\nu$ is defined as:
$\left\lVert\mu-\nu\right\rVert_{TV}=2\sup_{A\in\mathcal{B}}\left|\mu(A)-\nu(A)\right|=\int_{\Omega}\left|d\mu-d\nu\right|.$ (43)
If $\mu$ and $\nu$ have well defined densities $g,h$ with respect to an underlying measure $\rho$, then:
[…]
Below we give some equivalent definitions of the total variation distance:
Proposition B.2.
  • (i)
    We have:
[…]
  • (ii)
    Let $Y,Z$ be a coupling of $\mu$ and $\nu$, i.e. the marginal distribution of $Y$ is $\mu$ and the marginal distribution of $Z$ is $\nu$. Then:
    $\left\lVert\mu-\nu\right\rVert_{TV}=2\inf\left\{P\left(Y\neq Z\right),\text{$\left(Y,Z\right)$ is a coupling of $\mu,\nu$}\right\}.$ (46)
    The coupling $(Y,Z)$ that satisfies $\left\lVert\mu-\nu\right\rVert_{TV}=P\left(Y\neq Z\right)$ is then called the maximal coupling of $\mu$ and $\nu$.
The proof can be found in Proposition 3, [Roberts and Rosenthal, 2004].
Now we present some nice properties on the total variation distance w.r.t the stationary distribution of Markov chains:
Proposition B.3.
Let $X$ be a Markov chain on $\left(\Omega,\mathcal{B}\right)$ with transition kernel $P$ and stationary distribution $\pi$, then:
  • (i)
    For all probability distributions $\mu,\nu$ defined over $\left(\Omega,\mathcal{B}\right)$:
    $\left\lVert\mu P-\nu P\right\rVert_{TV}\leq\left\lVert\mu-\nu\right\rVert_{TV}.$ (47)
  • (ii)
    More specifically, advancing the chain will not increase the total variation distance to the stationary distribution, i.e.:
    $\left\lVert P^{n}\left(x,.\right)-\pi(.)\right\rVert_{TV}\leq\left\lVert P^{n-1}\left(x,.\right)-\pi(.)\right\rVert_{TV}.$ (48)
  • (iii)
    Let $d(n)=\sup_{x\in X}\left\lVert P^{n}\left(x,.\right)-\pi(.)\right\rVert_{TV}$. Then for $\mu$ varying over probability distribution defined over $\left(\Omega,\mathcal{B}\right)$:
    $d(n)=\sup_{\mu}\left\lVert\mu P^{n}-\pi\right\rVert_{TV}.$ (49)
    and $d(.)$ is sub-multiplicative, i.e.:
    $d(m+n)\leq d(m)d(n)\quad\text{for $m,n\in\mathbb{N}$.}$ (50)
  • (iv)
    Let $\left\{X_{0},X_{1},\dots,X_{n}\right\}$ be a finite sequence from the chain. Let $\mu,\nu$ two probability distribution defined over $\left(\Omega,\mathcal{B}\right)$, and $\mathcal{L}_{\mu},\mathcal{L}_{\nu}$ be the joint distribution of $\left\{X_{1},X_{2},\dots,X_{n}\right\}$ with $X_{1}\sim\mu$ and $X_{1}\sim\nu$ respectively. Then we have:
    $\left\lVert\mathcal{L}_{\mu}-\mathcal{L}_{\nu}\right\rVert_{TV}=\left\lVert\mu-\nu\right\rVert_{TV}.$ (51)
  • (v)
    For $i\in[n]$, let $\mu_{i}$ and $\nu_{i}$ be probability distributions on $\left(\Omega_{i},\mathcal{B}_{i}\right)$. Define $\mu=\bigotimes_{i=1}^{n}\mu_{i}$ and $\nu=\bigotimes_{i=1}^{n}\nu_{i}$ as two probability distributions on $(\Omega:=\bigtimes_{i=1}^{n}\Omega_{i},\mathcal{B}:=\bigotimes_{i=1}^{n}\mathcal{B}_{n})$. Then we have:
    $\left\lVert\mu-\nu\right\rVert_{TV}\leq\sum_{i=1}^{n}\left\lVert\mu_{i}-\nu_{i}\right\rVert_{TV}.$ (52)
Proof.
The proof of $(i),(ii)$ and $(iii)$ can also be found in Proposition 3, [Roberts and Rosenthal, 2004]. Here we present the proof for the last two statements.
  • (iv)
    Consider the "cylinder set" $A=A_{0}\times A_{1}\times\dots A_{n}\in\bigotimes_{i=0}^{n}\mathcal{B}$. Then by the Markov property, we have:
    $\begin{split}\mathcal{L}_{\mu}\left(A\right)&=\int_{A_{0}}\int_{A_{1}}\dots\int_{A_{n-1}}\mu\left(dx_{0}\right)P(x_{0},dx_{1})\dots P\left(x_{n-1},A_{n}\right),\\ \mathcal{L}_{\nu}\left(A\right)&=\int_{A_{0}}\int_{A_{1}}\dots\int_{A_{n-1}}\nu\left(dx_{0}\right)P(x_{0},dx_{1})\dots P\left(x_{n-1},A_{n}\right).\end{split}$ (53)
    Hence,
    $\begin{split}&\left\lVert\mathcal{L}_{\mu}-\mathcal{L}_{\nu}\right\rVert_{TV}\\ &=2\sup_{A_{0}}\left|\int_{A_{0}}\mu\left(dx_{0}\right)-\nu\left(dx_{0}\right)\right|\sup_{A_{1},\dots A_{n}}\left|\int_{A_{1}}\dots\int_{A_{n-1}}P\left(x_{0},dx_{1}\right)\dots P\left(x_{n-1},A_{n}\right)\right|\\ &=\left\lVert\mu-\nu\right\rVert_{TV}\left|\int_{\Omega}\dots\int_{\Omega}P(x_{0},dx_{1})\dots P(x_{n-1},\Omega)\right|\\ &=\left\lVert\mu-\nu\right\rVert_{TV}.\end{split}$ (54)
  • (v)
    By using (46), it follows that
    $\left\lVert\mu_{m}-\nu_{m}\right\rVert_{TV}=\mathbb{P}_{m}\left\{X_{m}\neq Y_{m}\right\},$ (55)
    where $\left(X_{m},Y_{m}\right)$ is the maximal coupling of $\mu_{m},\nu_{m}$. Let’s define $X=\left\{X_{m}\right\}_{m\in[M]}$ and $Y=\left\{Y_{m}\right\}_{m\in[M]}$. Then again by (46) we have
    $\begin{split}\left\lVert\mu-\nu\right\rVert_{TV}&\leq\mathbb{P}\left\{X\neq Y\right\}\\ &=\mathbb{P}\left\{\bigcup_{m=1}^{M}X_{m}\neq Y_{m}\right\}\\ &\leq\sum_{m=1}^{M}\mathbb{P}_{m}\left\{X_{m}\neq Y_{m}\right\}=\sum_{m=1}^{M}\left\lVert\mu_{m}-\nu_{m}\right\rVert_{TV},\end{split}$ (56)
    which follows from the union bound.
Proposition B.4.
Given a Markov chain defined on $\left(\Omega,\mathcal{B}\right)$, let $d(t)$ the total variation distance defined in (49), $\tau\left(\epsilon\right)$ and $\tau$ defined in Definition 3.3, for any positive integer $n$, we have
$d(n\tau\left(\epsilon\right))\leq\epsilon^{n}.$ (57)
More specifically,
$d\left(n\tau\right)\leq 4^{-n},$ (58)
and
$\tau\left(\epsilon\right)\leq\lceil\log_{4}\left(\epsilon^{-1}\right)\rceil\tau.$ (59)
Proof.
The first statement is a direct application of (50). Simply by taking $\epsilon=1/4$, we obtain the second statement. Then with $n\geq\lceil\log_{4}\left(\epsilon^{-1}\right)\rceil$, $d\left(n\tau\right)\leq\epsilon$, which implies that $n\tau\geq\tau\left(\epsilon\right)$, hence the last statement. ∎
Proposition B.5.
Given $M$ independent time-homogeneous Markov chain, each defined on $\left(\Omega_{m},\mathcal{B}_{m}\right)$, with transition kernel $P_{m}$, stationary distribution $\pi_{m}$ and finite mixing time $\tau_{m}$, where $m\in[M]$. Consider the product Markov chain defined on $(\Omega:=\bigtimes_{m=1}^{M}\Omega_{m},\mathcal{B}:=\bigotimes_{m=1}^{M}\mathcal{B}_{m})$ with transition kernel given by the following tensor product:
$P\left(\left(x_{1},\dots,x_{M}\right),\left(A_{1},\dots,A_{M}\right)\right)=\prod_{m=1}^{M}P_{m}\left(x_{m},A_{m}\right)\quad\textrm{with $x_{m}\in\Omega_{m},A_{m}\in\mathcal{B}_{m}$, for all $m\in[M]$}.$ (60)
The stationary distribution of this product chain is $\pi=\bigotimes_{m=1}^{M}\pi_{m}$. Then the mixing time of this product Markov chain $\tau$ satisfies
$\tau\leq\left(\lceil\log_{4}M\rceil+1\right)\underset{m\in[M]}{\max}\tau_{m}.$ (61)
Proof.
Let
$\begin{split}d(t)&=\underset{x\in\mathcal{X}}{\sup}\left\lVert P^{t}\left(x,.\right)-\pi\right\rVert,\\ d_{m}(t)&=\underset{x\in\mathcal{X}_{m}}{\sup}\left\lVert P_{m}^{t}\left(x,.\right)-\pi_{m}\right\rVert.\end{split}$ (62)
By Proposition B.3, we have
$d(t)\leq\sum_{m=1}^{M}d_{m}(t).$ (63)
As in (48), advancing the Markov chain in time can only move it closer to the stationary distribution. Thus, for all $m\in[M]$,
$d_{m}\left(\underset{m\in[M]}{\max}\tau_{m}\left(\frac{\epsilon}{M}\right)\right)\leq\frac{\epsilon}{M},$ (64)
where $\tau\left(\epsilon\right):=\min\left\{t:d(t)\leq\epsilon\right\}$ and $\tau_{m}\left(\epsilon\right)=\min\left\{t:d_{m}(t)\leq\epsilon\right\}$. Hence, from (63),
$\tau\left(\epsilon\right)\leq\underset{m\in[M]}{\max}\tau_{m}\left(\frac{\epsilon}{M}\right).$ (65)
Then from Proposition B.4, we have that, for all $m\in[M]$,
$\tau_{m}\left(\epsilon/M\right)\leq\lceil\log_{4}\left(M/\epsilon\right)\rceil\tau_{m}.$ (66)
The lemma follows by taking $\epsilon=1/4$ as in the definition of mixing time (Definition 3.3). ∎
Appendix C Some useful lemmas
Lemma C.1.
Let $\mu$ and $\nu$ the joint probability distribution of $\left(x_{t}^{(m,k)}\right)_{m\in[M],k\in[K]}$ with initial distribution $\mathcal{Q}$ of $\left(x_{0}^{(m,0)}\right)_{m\in[M]}$ and $\pi=\bigotimes_{m=1}^{M}\pi_{m}$ respectively. Assume that $\mu$ and $\nu$ has well-defined densities $g$ and $h$ with respect to a measure $\rho$ on $\left(\Omega,\mathcal{B}\right)$ as defined in Section 3.2. If Assumptions 3.13.4 and 4.3 hold, then
$\mathbb{E}_{\mathcal{Q}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\tau\sigma^{2}}{MK}+\sigma^{2}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.$ (67)
where $P_{\mathcal{Q}}$ and $\mathbb{E}_{\mathcal{Q}}$ denote the data distribution and expectation, respectively, with initial state distribution $\mathcal{Q}$.
Proof.
We have
$\begin{split}\mathbb{E}_{\mathcal{Q}}&\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla F_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ =&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\int_{\Omega}\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\left(g\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)-h\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\right)\mathrm{d}\rho\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\\ \overset{(i)}{\leq}&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\sigma^{2}\int_{\Omega}\left\lvert g\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)-h\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\right\rvert\mathrm{d}\rho\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\\ \overset{(ii)}{\leq}&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\sigma^{2}\left\lVert\mu-\nu\right\rVert_{TV}.\end{split}$ (68)
where in $(i)$ we use the uniform bound on $\left\lVert\nabla F_{m}\left(w;x\right)-\nabla F_{m}\left(w\right)\right\rVert^{2}$ (Assumption 4.3) and in $(ii)$, we use the definition of the total variation distance (Definition B.1).
To bound the first term, we utilize [Paulin, 2015, Theorem 3.2], which provides a concentration inequality for the case of stationary Markov chains.This yields
$\begin{split}\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq&\frac{1}{\gamma_{ps}K}\mathbb{E}_{\pi}\left[\left\lVert\frac{1}{M}\sum_{m}\nabla f_{m}\left(w_{t};x\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ =&\frac{1}{M^{2}K\gamma_{ps}}\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right].\\ \end{split}$ (69)
Here $\gamma_{ps}$ is the pseudo spectral gap of the transition kernel $P$ of the system-level Markov chain (see [Paulin, 2015, Section 3.1]). With Assumption 3.4[Paulin, 2015, Proposition 3.4] gives $\frac{1}{\gamma_{ps}}\leq 2\tau$. Hence we have:
$\begin{split}\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq&\frac{2\tau}{M^{2}K}\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right]\\ \leq&\frac{2\tau\sigma^{2}}{MK}.\end{split}$ (70)
The last step follows from the fact $\mathbb{E}_{\pi_{m}}\left[\nabla f_{m}\left(w;x\right)\right]=\nabla F_{m}\left(w\right)$. As such, the term $\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]$ can be seen as the variance of the sum of $M$ independent variables. Using Assumption 4.3, it then follows that $\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right]=\sum_{m}\mathbb{E}_{\pi_{m}}\left[\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right]\leq M\sigma^{2}$.
Using (51), we also have
$\begin{split}\left\lVert\mu-\nu\right\rVert_{TV}&=\left\lVert P_{\mathcal{Q}}\left(\left(x_{t}^{(m,0)}\right)_{m\in[M]},.\right)-\pi\right\rVert_{TV}\\ &=\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (71)
Hence we prove the lemma. ∎
Lemma C.2.
Given Assumption 3.4, with $K\geq\tau$, we have that:
$\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\leq\frac{2\tau}{K\ln 4}+1.$ (72)
Proof.
Since we assume that the product chain is uniformly ergodic, we have that:
$\begin{split}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\leq\left(\frac{1}{4}\right)^{\lfloor Kt/\tau\rfloor}=\left(\frac{1}{4}\right)^{\lfloor K/\tau\rfloor t}.\end{split}$ (73)
By using the geometric sum formula, we have:
$\begin{split}\sum_{t=0}^{T-1}\left(\frac{1}{4}\right)^{\lfloor K/\tau\rfloor t}&\leq\frac{1}{1-(1/4)^{\lfloor K/\tau\rfloor}}\\ &=1+\frac{1}{4^{\lfloor K/\tau\rfloor}-1}\\ &=1+\frac{1}{\exp\left(\lfloor K/\tau\rfloor\ln 4\right)-1}\\ &\leq\frac{1}{\lfloor K/\tau\rfloor\ln 4}+1.\end{split}$ (74)
Here in the last step, we use the inequality $e^{x}\geq 1+x$ for all $x$. Since $K\geq\tau$, we have that $\lfloor K/\tau\rfloor\geq K/(2\tau)$, which proves the lemma. ∎
Appendix D Minibatch SGD
In this section, we consider the problem class $\mathcal{F}_{1}\left(L,\sigma,\tau\right)$. We begin with the following descent lemma.
Lemma D.1.
For the problem class $\mathcal{F}_{1}(L,\sigma,\tau)$, with constant step size $\gamma\leq\frac{1}{L}$, the iterates of Minibatch SGD satisfy:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{\gamma\tau\sigma^{2}}{MK}+\frac{\gamma\sigma^{2}}{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (75)
Proof.
Since $F$ is $L$-smooth,
$\begin{split}F\left(w_{t+1}\right)&\leq F\left(w_{t}\right)+\langle\nabla F\left(w_{t}\right),w_{t+1}-w_{t}\rangle+\frac{L}{2}\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\\ &=F\left(w_{t}\right)-\gamma\langle\nabla F\left(w_{t}\right),\bar{g}_{t}\rangle+\frac{L}{2}\left\lVert w_{t+1}-w_{t}\right\rVert^{2}.\end{split}$ (76)
Note that the term $-\gamma\langle\nabla F\left(w_{t}\right),\bar{g}_{t}\rangle$ can be expanded as
$-\gamma\langle\nabla F\left(w_{t}\right),\bar{g}_{t}\rangle$ $=-\frac{\gamma}{2}\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}+\frac{\gamma}{2}\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}-\frac{\gamma}{2}\left\lVert\bar{g}_{t}\right\rVert^{2}$ (77)
$=-\frac{\gamma}{2}\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}+\frac{\gamma}{2}\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}-\frac{1}{2\gamma}\left\lVert w_{t+1}-w_{t}\right\rVert^{2}.$ (78)
Plugging (77) back into (76), subtracting $F^{*}$ from both sides, and taking expectation we then have
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]-\left(\frac{1}{2\gamma}-\frac{L}{2}\right)s_{t}.$ (79)
Then with the condition $\gamma\leq\frac{1}{L}$, the last term in (79) can be ignored. The lemma follows by replacing the second term in (79) using Lemma C.1. ∎
Now we proceed to the proof of Theorem 4.5.
Theorem D.2 (Theorem 4.5 in the main paper).
For the problem class $\mathcal{F}_{1}(L,\sigma,\tau)$, with global step size $\gamma\leq 1/L$, the iterates of Minibatch SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\gamma T}+\frac{\tau\sigma^{2}}{MK}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (80)
Proof.
The proof of Theorem D.2 follows directly from Lemma D.1 by rearranging the terms and taking the average from $t=0$ to $T-1$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{2\tau\sigma^{2}}{MK}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (81)
Noticing that with $\hat{w}_{T}$ sampled uniformly from $w_{0},\dots w_{T-1}$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]=\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right],$
which concludes the proof. ∎
Corollary 4.6 can be derived directly from Theorem D.2.
Corollary D.3 (Corollary 4.6 in the main paper).
Under the conditions of Theorem D.2, to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$, the required number of local steps and communication rounds of Minibatch SGD are:
  1. (i)
    The system-level Markov chain $X$ is stationary $\left(Q=\pi\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right).$ (82)
  2. (ii)
    The system-level Markov chain $X$ is non-stationary $\left(Q\neq\pi\right)$, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (83)
Proof.
In the case where the system-level Markov chain is stationary, since $\pi P=\pi$, the term $\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}$ in Theorem D.2 disappears. Choosing
$\begin{split}K&\geq\frac{4\tau\sigma^{2}}{M\epsilon^{2}},\\ T&\geq\frac{4\Delta_{0}}{\gamma\epsilon^{2}}\geq\frac{4L\Delta_{0}}{\epsilon^{2}},\end{split}$
guarantees that $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$.
In the case where the system-level Markov chain is non-stationary, by applying Lemma C.2 to Theorem D.2, we have
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{2\tau\sigma^{2}}{MK}+\frac{2\tau\sigma^{2}}{TK\ln 4}+\frac{\sigma^{2}}{T}.$
Choosing
$\begin{split}K&\geq\frac{8\tau\sigma^{2}}{M\epsilon^{2}},\\ T&\geq\frac{8\left(L\Delta_{0}\vee\sigma^{2}\right)}{\epsilon^{2}},\end{split}$
with
$\epsilon\leq\sqrt{\frac{8\sigma^{2}}{M}}$ (84)
to ensure $K\geq\tau$. This yield:
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{3}{4}\epsilon^{2}+\frac{M\epsilon^{2}}{64\ln 4\left(L\Delta_{0}\vee\sigma^{2}\right)}\epsilon^{2}.$
Hence, in the high-precision regime with
$\epsilon\leq\sqrt{\frac{8\sigma^{2}}{M}}\leq\sqrt{\frac{16\ln 4\left(L\Delta_{0}\vee\sigma^{2}\right)}{M}},$
we conclude that $\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\epsilon^{2}$. ∎
Appendix E Local SGD
For the analysis of Local SGD, we define
$\xi_{t}=\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right],$
where $\xi_{t}$ is the client drift observed in the local computation phase of the $t$-th communication round.
We recall that the analysis of Local SGD is for the problem class $\mathcal{F}_{3}\left(L,\sigma,\theta,\delta,\tau\right)$. We begin with the following descent lemma.
Lemma E.1.
If the constant step size satisfies
$\tilde{\eta}\leq\frac{1}{10L\delta^{2}}<\frac{1}{L},$ (85)
then the iterates of Local SGD satisfy
$\Delta_{t+1}\leq\Delta_{t}-\frac{45\tilde{\eta}}{98}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{2\tilde{\eta}\tau\sigma^{2}}{MK}+\frac{40}{98}\frac{L\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\tilde{\eta}\sigma^{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (86)
Proof.
We can adapt the proof of Lemma D.1 to Local SGD but with the update rule $w_{t+1}=w_{t}-\tilde{\eta}g_{t}$:
$\begin{split}\Delta_{t+1}&\leq\Delta_{t}-\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert g_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]-\left(\frac{1}{2\tilde{\eta}}-\frac{L}{2}\right)s_{t}\\ &\leq\Delta_{t}-\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\tilde{\eta}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\tilde{\eta}\mathbb{E}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right].\end{split}$ (87)
Here in the last step, we use the triangle inequality, and the condition on the step size $\tilde{\eta}\leq\frac{1}{10L\delta^{2}}<\frac{1}{L}$.
Using the sample-wise smoothness Assumption 4.2 yields
$\begin{split}\mathbb{E}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right]&=\mathbb{E}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)\right\rVert^{2}\right]\\ &\leq\frac{L^{2}}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ &=L^{2}\xi_{t}.\end{split}$ (88)
where the triangle inequality is used in the last step.
The local drift $\xi_{t}$ can then be bounded as follows:
$\begin{split}\xi_{t}=&\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ =&\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert\sum_{j=1}^{k}\eta\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{1}{MK}\sum_{m,k}\eta^{2}k\sum_{j=1}^{k}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=1}^{k}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=1}^{K}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ &+\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \overset{(i)}{\leq}&\quad 2L^{2}\eta^{2}\xi_{t}+\frac{4\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)-\nabla F_{m}(w_{t})\right\rVert^{2}\right]\\ &+\frac{4\tilde{\eta}^{2}}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\\ \overset{(ii)}{\leq}&2L^{2}\tilde{\eta}^{2}\xi_{t}+4\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)+4\tilde{\eta}^{2}\delta^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (89)
In $(i)$ we use the sample-wise smoothness (Assumption 4.2) and the triangle inequality, and in $(ii)$ we use the uniform bound on the noise of the gradient estimator (Assumption 4.3) and the heterogeneity assumption (Assumption 4.4).
Therefore we have:
$\begin{split}\xi_{t}\leq\frac{4\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)}{1-2L^{2}\tilde{\eta}^{2}}+\frac{4\tilde{\eta}^{2}\delta^{2}}{1-2L^{2}\tilde{\eta}^{2}}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (90)
Plugging (88) and (90) back into (87), together with Lemma C.1 and the condition on the step size then yields
$\begin{split}\Delta_{t+1}\leq&\Delta_{t}-\frac{45\tilde{\eta}}{98}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\frac{2\tilde{\eta}\tau\sigma^{2}}{MK}+\frac{40}{98}L\tilde{\eta}^{2}\frac{\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\tilde{\eta}\sigma^{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV},\end{split}$ (91)
which concludes the proof of the lemma. ∎
We now prove Theorem 4.7.
Theorem E.2 (Theorem 4.7 in the main paper).
For the problem class $\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, if the constant step size satisfies
$\eta\leq\mathcal{O}\left(\frac{1}{LK\delta^{2}}\right)<\frac{1}{L},$ (92)
then the iterates of Local SGD satisfy
$\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\eta KT}+\frac{\tau\sigma^{2}}{MK}+\frac{L\eta K\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).$ (93)
Proof.
By rearranging the terms in Lemma E.1 and taking the average from $t=0$ to $T-1$:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]&\leq\frac{98}{45}\frac{\Delta_{0}}{\tilde{\eta}T}+\frac{196}{45}\frac{\tau\sigma^{2}}{MK}+\frac{40}{45}\frac{L\tilde{\eta}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{98}{45}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.\\ \end{split}$ (94)
We conclude the proof by replacing $\tilde{\eta}=\eta K$, and noticing that with $\hat{w}_{T}$ sampled uniformly from $w_{0},\dots w_{T-1}$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]=\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right].$
We now establish Corollary 4.8, which characterizes the communication and sample complexities of Local SGD.
Corollary E.3 (Corollary 4.8 in the main paper).
For the problem class $\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, with step size
$\eta\leq\mathcal{O}\left(\frac{\delta^{2}\epsilon^{2}}{KL(\theta^{2}+\sigma^{2})}\wedge\frac{1}{KL\delta^{2}}\right),$ (95)
the required number of local steps and communication rounds for Local SGD to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are given by
  1. (i)
    If the system-level Markov chain $X$ is stationary $\left(Q=\pi\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (96)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (97)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary $\left(Q\neq\pi\right)$, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (98)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (99)
Proof.
In the case where the system-level Markov chain is stationary, since $\pi P=\pi$, the term $\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}$ disappears. Choosing
$\begin{split}K&\geq\frac{196}{15}\frac{\tau\sigma^{2}}{M\epsilon^{2}},\\ \eta&\leq\left(\frac{3}{8}\frac{\delta^{2}\epsilon^{2}}{KL\left(\theta^{2}+\sigma^{2}\right)}\wedge\frac{1}{10KL\delta^{2}}\right),\\ T&\geq\frac{98}{15}\frac{\Delta_{0}}{\eta K\epsilon^{2}}\geq\frac{98}{15}\frac{L\Delta_{0}}{\epsilon^{2}}\left(10\delta^{2}\vee\frac{8}{3}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right),\end{split}$
and applying Theorem E.2, then yields $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$.
In the case where the system-level Markov chain is non-stationary, by applying Lemma C.2 to Theorem E.2, we have
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{98}{45}\frac{\Delta_{0}}{\eta KT}+\frac{196}{45}\frac{\tau\sigma^{2}}{MK}+\frac{40}{45}\frac{L\eta K\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{98}{45}\frac{\sigma^{2}}{T}+\frac{196}{45}\frac{\tau\sigma^{2}}{KT\ln 4}.$
Choosing
$\begin{split}K&\geq\frac{196}{9}\frac{\tau\sigma^{2}}{M\epsilon^{2}},\\ \eta&\leq\left(\frac{9}{40}\frac{\delta^{2}\epsilon^{2}}{KL\left(\theta^{2}+\sigma^{2}\right)}\wedge\frac{1}{10KL\delta^{2}}\right),\\ T&\geq\frac{98}{9}\frac{\Delta_{0}\vee\sigma^{2}}{\eta K\epsilon^{2}}\geq\frac{98}{9}\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right),\end{split}$
with
$\epsilon\leq\sqrt{\frac{196}{9}\frac{\sigma^{2}}{M}}$ (100)
to ensure that $K\geq\tau$. It follows that
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{4}{5}\epsilon^{2}+\frac{9}{490\ln 4}\frac{M\epsilon^{4}}{\left(L\Delta_{0}\vee\sigma^{2}\right)\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)}.$
In the high-precision regime,
$\epsilon\leq\sqrt{\frac{98\ln 4}{9}\frac{\sigma^{2}}{M}}\leq\sqrt{\frac{98\ln 4}{9}\frac{\sigma^{2}}{M}\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)},$
it then follows that $\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\epsilon^{2}$. ∎
Appendix F Local SGD with Momentum
All the proofs in this section are for the problem class $\mathcal{F}_{2}\left(L,\sigma,\tau\right)$. Recall that the update rule of Local SGD-M can be written as: $w_{t+1}=w_{t}-\gamma v_{t+1}$. For the theoretical analysis of Local SGD with momentum, we introduce the following notation:
$\begin{split}\Xi_{t}&=\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)-v_{t+1}\right\rVert^{2}\right],\\ G_{0}&=\frac{1}{M}\sum_{m}\left\lVert\nabla F\left(w_{0}\right)\right\rVert^{2}.\end{split}$ (101)
We begin with the following descent lemma.
Lemma F.1.
With global step size
$\gamma\leq\frac{1}{L},$ (102)
we have
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\gamma}{2}\Xi_{t}.$ (103)
Proof.
Similarly as the proof of Lemma D.1, with the update rule of Local SGD-M: $w_{t+1}=w_{t}-\gamma v_{t+1}$, we have:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\gamma}{2}\Xi_{t}+\left(\frac{L}{2}-\frac{\gamma}{2}\right)s_{t}.$ (104)
Applying the condition on $\gamma$ proves the lemma. ∎
We have the following recursive bound on $\Xi_{t}$:
Lemma F.2.
With condition:
$\gamma\leq\frac{1}{\sqrt{60}}\frac{\beta}{L}.$ (105)
We have:
$\Xi_{t}\leq\left(1-\frac{14}{15}\beta\right)\Xi_{t-1}+\frac{\beta}{15}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]+6\beta\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+6\beta L^{2}\xi_{t}.$ (106)
Proof.
We have the following decomposition:
$\begin{split}\nabla F\left(w_{t}\right)-v_{t+1}&=\left(1-\beta\right)\left(\nabla F\left(w_{t}\right)-v_{t}\right)-\beta\left(\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right)\\ &-\frac{\beta}{MK}\sum_{m,k}\left(\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)\right)\\ &=\left(1-\beta\right)\left(\nabla F\left(w_{t}\right)-v_{t}\right)-\beta\left(\bar{g}_{t}-\nabla F(w_{t})\right)-\beta\left(g_{t}-\bar{g}_{t}\right).\end{split}$
Hence, by using Young inequality:
$\left(a+b\right)^{2}\leq\left(1+\frac{\beta}{2}\right)a^{2}+\left(1+\frac{2}{\beta}\right)b^{2}.$
We have:
$\begin{split}\Xi_{t}\leq&\left(1+\frac{\beta}{2}\right)\left(1-\beta\right)^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)-v_{t}\right\rVert^{2}\right]\\ &+\left(1+\frac{2}{\beta}\right)\beta^{2}\mathbb{E}\left[\left\lVert\left(g_{t}-\bar{g}_{t}\right)-\left(\bar{g}_{t}-\nabla F\left(w_{t}\right)\right)\right\rVert^{2}\right]\\ \leq&\left(1+\frac{\beta}{2}\right)\left(1-\beta\right)^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)-v_{t}\right\rVert^{2}\right]+2\left(1+\frac{2}{\beta}\right)\beta^{2}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+2\left(1+\frac{2}{\beta}\right)\beta^{2}L^{2}\xi_{t}.\end{split}$ (107)
Here in the last step, we use triangle inequality and the sample-wise smoothness (Assumption 4.2).
Now, since we have:
$\begin{split}\nabla F(w_{t})-v_{t}&=\nabla F(w_{t-1})-v_{t}+\left(\nabla F(w_{t})-\nabla F(w_{t-1})\right)\\ &=\Xi_{t-1}+\left(\nabla F(w_{t})-\nabla F(w_{t-1})\right).\end{split}$ (108)
Using again Young inequality we have:
$\begin{split}\Xi_{t}\leq&(1+\frac{\beta}{2})\left(1-\beta\right)^{2}\left[\left(1+\frac{\beta}{2}\right)\Xi_{t-1}+\left(1+\frac{2}{\beta}\right)\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)-\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]\right]\\ &+6\beta\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+6\beta L^{2}\xi_{t}\\ \leq&\left(1-\beta\right)\Xi_{t-1}+\frac{2L^{2}}{\beta}\mathbb{E}\left[\left\lVert w_{t}-w_{t-1}\right\rVert^{2}\right]+6\beta\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+6\beta L^{2}\xi_{t}.\end{split}$ (109)
Here since $\beta\leq 1$, in the first inequality we use $1+\frac{2}{\beta}\leq\frac{3}{\beta}$; in the second inequality we use $\left(1+\frac{\beta}{2}\right)^{2}\left(1-\beta\right)^{2}\leq\left(1-\beta\right)$ and $\left(1+\frac{\beta}{2}\right)\left(1+\frac{2}{\beta}\right)\left(1-\beta\right)\leq\frac{2}{\beta}$ together with the smoothness of $F$.
Furthermore, by the update rule, we have:
$\begin{split}\mathbb{E}\left[\left\lVert w_{t}-w_{t-1}\right\rVert^{2}\right]=&\quad\gamma^{2}\mathbb{E}\left[\left\lVert v_{t-1}\right\rVert^{2}\right]\\ \leq&\quad 2\gamma^{2}\Xi_{t-1}+2\gamma^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right].\end{split}$ (110)
Plugging (110) back into (109) we have:
$\begin{split}\Xi_{t}\leq&\left(1-\beta+\frac{4\gamma^{2}L^{2}}{\beta}\right)\Xi_{t-1}+\frac{4\gamma^{2}L^{2}}{\beta}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]+6\beta\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+6\beta L^{2}\xi_{t}.\end{split}$ (111)
We conclude the proof by using the condition on the global step size $\gamma$. ∎
Next, we continue to bound the drift $\xi_{t}$.
Lemma F.3.
With the following conditions:
$\eta KL\leq\frac{1}{2\beta}.$ (112)
We have:
$\begin{split}\xi_{t}\leq&\quad 16\left(\eta\beta K\right)^{2}\sigma^{2}+32\left[\eta\left(1-\beta\right)K\right]^{2}\Xi_{t-1}+32\left[\eta\left(1-\beta\right)K\right]^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]\\ &+16\left(\eta\beta K\right)^{2}\frac{1}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (113)
Proof.
We start by using Young inequality $\left(a+b\right)^{2}\leq\left(1+\frac{1}{K-1}\right)a^{2}+Kb^{2}$. Then, for any $10\leq k\leq K-$ 1 we have:
$\begin{split}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\leq&\quad\left(1+\frac{1}{K-1}\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]+K\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}^{(m,k)}\right\rVert^{2}\right]\\ =&\quad\left(1+\frac{1}{K-1}\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]\\ &+K\eta^{2}\mathbb{E}\left[\left\lVert\beta\nabla f_{m}\left(w_{t}^{(m,k-1)};x_{t}^{(m,k-1)}\right)+\left(1-\beta\right)v_{t}\right\rVert^{2}\right].\\ \end{split}$ (114)
Now we have the following decomposition:
$\begin{split}\beta\nabla f_{m}\left(w_{t}^{(m,k-1)};x_{t}^{(m,k-1)}\right)+\left(1-\beta\right)v_{t}=&\left(1-\beta\right)v_{t}+\beta\nabla F_{m}\left(w_{t}\right)\\ &+\beta\left[\nabla f_{m}\left(w_{t}^{(m,k-1)};x_{t}^{(m,k-1)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k-1)}\right)\right]\\ &+\beta\left[\nabla f_{m}\left(w_{t};x_{t}^{(m,k-1)}\right)-\nabla F_{m}\left(w_{t}\right)\right].\end{split}$ (115)
Hence, by using triangle inequality we have:
$\begin{split}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\leq&\left(1+\frac{1}{K-1}\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]\\ &+4\eta^{2}\beta^{2}K\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,k-1)};x_{t}^{(m,k-1)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k-1)}\right)\right\rVert^{2}\right]\\ &+4\eta^{2}\beta^{2}K\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,k-1)}\right)-\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\\ &+4\eta^{2}\beta^{2}\mathbb{E}\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]\\ &+8\eta^{2}\left(1-\beta\right)^{2}K\left(\Xi_{t-1}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]\right)\\ \leq&\left(1+\frac{1}{K-1}+4\eta^{2}\beta^{2}L^{2}K\right)\mathbb{E}\left[\left\lVert w_{t}^{(m,k-1)}-w_{t}\right\rVert^{2}\right]+4K\mathcal{A},\end{split}$ (116)
with
$\mathcal{A}=\eta^{2}\beta^{2}\left(\sigma^{2}+\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\right)+2\eta^{2}\left(1-\beta\right)^{2}\left(\Xi_{t-1}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]\right).$ (117)
Here we use the sample-wise smoothness and the uniform bound on the gradient noise in the last step (Assumptions 4.24.3).
Unrolling the above recursion, with $w_{t}^{(m,0)}=w_{t}$, we have:
$\begin{split}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\leq&\quad 4K\mathcal{A}\sum_{j=0}^{k-1}\left(1+\frac{1}{K-1}+4\eta^{2}\beta^{2}L^{2}K\right)^{j}.\end{split}$ (118)
With the condition on $\eta$, we have that $4\eta^{2}\beta^{2}L^{2}K\leq\frac{1}{K}<\frac{1}{K-1}$. Hence by using the geometric sum formula, we have:
$\begin{split}\sum_{j=0}^{k-1}\left(1+\frac{1}{K-1}+4\eta^{2}\beta^{2}L^{2}K\right)^{j}\leq&\frac{1-\left(1+\frac{2}{K-1}\right)^{k}}{-\frac{2}{K-1}}\\ =&\frac{K-1}{2}\left[\left(1+\frac{2}{K-1}\right)^{k}-1\right]\\ \leq&\frac{K-1}{2}\left(1+\frac{2}{K-1}\right)^{K-1}\\ \leq&\frac{e^{2}}{2}\left(K-1\right)\leq 4K.\end{split}$ (119)
We conclude the lemma by taking the average over $K$ and $M$. ∎
In the proof of Local SGD we have to use the heterogeneity assumption to bound the term $\frac{1}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]$. The use of momentum allows us to build a recursive bound for this term by the following lemma.
Lemma F.4.
We have that:
$\frac{1}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\leq 3G_{0}+6\left(1+t\right)\gamma^{2}L^{2}\sum_{j=0}^{t-1}\left(\Xi_{j}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{j}\right)\right\rVert^{2}\right]\right).$ (120)
Proof.
By using Young inequality and the smoothness of $F_{m}$, we have that:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\leq&\left(1+a\right)\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t-1}\right)\right\rVert^{2}\right]+\left(1+a^{-1}\right)L^{2}\mathbb{E}\left[\left\lVert w_{t}-w_{t-1}\right\rVert^{2}\right]\\ \leq&\left(1+a\right)\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t-1}\right)\right\rVert^{2}\right]+2\left(1+a^{-1}\right)\gamma^{2}L^{2}\left(\Xi_{t-1}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{t-1}\right)\right\rVert^{2}\right]\right).\end{split}$ (121)
By unrolling we have:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\leq&\left(1+a\right)^{t}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{0}\right)\right\rVert^{2}\right]+2\left(1+a^{-1}\right)\gamma^{2}L^{2}\sum_{j=0}^{t-1}\left(\Xi_{j}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{j}\right)\right\rVert^{2}\right]\right)\left(1+a\right)^{t-1-j}\\ \leq&e^{at}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{0}\right)\right\rVert^{2}\right]+2e^{at}\left(1+a^{-1}\right)\gamma^{2}L^{2}\sum_{j=0}^{t-1}\left(\Xi_{j}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{j}\right)\right\rVert^{2}\right]\right).\end{split}$ (122)
By taking $a=t^{-1}$ and taking the average over $m\in[M]$, we conclude the lemma. ∎
Lemma F.5.
With the following conditions:
$\begin{split}\eta KL\leq&\left(\frac{1}{2\beta}\wedge\frac{1}{\sqrt{120MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{360\beta^{3}TG_{0}}}\wedge\frac{1}{225}\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{525}\frac{1}{\beta\gamma LT}\right),\\ \gamma\leq&\frac{1}{\sqrt{60}}\frac{\beta}{L}.\end{split}$ (123)
We have that:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{7}{2}\frac{L\Delta_{0}}{\beta T}+\frac{17}{2}\frac{\tau\sigma^{2}}{MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{15}{2}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (124)
Proof.
By Lemma F.4, we have that:
$\begin{split}\frac{1}{M}\sum_{t=0}^{T-1}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\leq&3TG_{0}+6\gamma^{2}L^{2}\sum_{t=0}^{T-1}(1+t)\sum_{j=0}^{t-1}\left(\Xi_{j}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{j}\right)\right\rVert^{2}\right]\right)\\ \leq&3TG_{0}+12\left(\gamma LT\right)^{2}\sum_{t=0}^{T-2}\left(\Xi_{t}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\right).\end{split}$ (125)
Plugging (125) back into Lemma F.3 we have:
$\begin{split}6\beta L^{2}\sum_{t=0}^{T-1}\xi_{t}\leq&96\beta T\left(\eta KL\beta\right)^{2}\left(\sigma^{2}+3G_{0}\right)\\ &+\beta\underbrace{\left[192\left(\eta KL\left(1-\beta\right)\right)^{2}+1152\left(\eta KL\beta\right)^{2}\left(\gamma LT\right)^{2}\right]}_{\mathcal{B}}\sum_{t=0}^{T-2}\left(\Xi_{t}+\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\right).\end{split}$ (126)
By condition on the step size, we have that $\mathcal{B}\leq\frac{2}{15}$. Therefore, plugging (126) back into Lemma F.2, by noticing that $\mathbb{E}\left[\left\lVert\nabla F\left(w_{-1}\right)\right\rVert^{2}\right]=0$, gives:
$\begin{split}\sum_{t=0}^{T-1}\Xi_{t}\leq&\left(1-\frac{14}{15}\beta\right)\sum_{t=-1}^{T-2}\Xi_{t}+\frac{\beta}{15}\sum_{t=0}^{T-2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+6\beta\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F(w_{t})\right\rVert\right]+96\beta T\left(\eta KL\beta\right)^{2}\left(\sigma^{2}+3G_{0}\right)\\ &+\frac{2\beta}{15}\left(\sum_{t-1}^{T-2}\Xi_{t}+\sum_{t=0}^{T-2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\right)\\ \leq&\left(1-\frac{4}{5}\beta\right)\sum_{t=-1}^{T-2}\Xi_{t}+\frac{\beta}{5}\sum_{t=0}^{T-2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+6\beta\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F(w_{t})\right\rVert^{2}\right]+96\beta T\left(\eta KL\beta\right)^{2}\left(\sigma^{2}+3G_{0}\right).\end{split}$ (127)
By rearranging the terms, we have:
$\begin{split}\sum_{t=0}^{T-1}\Xi_{t}\leq&\quad\frac{5}{4\beta}\Xi_{-1}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{15}{2}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+120T\left(\eta KL\beta\right)^{2}\left(\sigma^{2}+3G_{0}\right).\end{split}$ (128)
With the condition on the step size, we have:
$\begin{split}120(\eta KL\beta)^{2}\sigma^{2}\leq&\frac{\sigma^{2}}{MK}\leq\frac{\tau\sigma^{2}}{MK}\quad\textrm{(since $\tau$ is always bigger than or equal to $1$)},\\ 360\left(\eta KL\beta\right)^{2}G_{0}\leq&\frac{L\Delta_{0}}{\beta T}.\end{split}$ (129)
Hence, by plugging (129)) and Lemma C.1 back into (128) we have:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}\leq\frac{5}{4}\frac{\Xi_{-1}}{\beta T}+\frac{L\Delta_{0}}{\beta T}+\frac{16\tau\sigma^{2}}{MK}+\frac{1}{4}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{15}{2}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (130)
Finally noticing that if we initialize $v_{0}=0$, $\Xi_{-1}=\mathbb{E}\left[\left\lVert\nabla F\left(w_{0}\right)\right\rVert^{2}\right]\leq 2L\Delta_{0}$ by the smoothness of $F$. Hence we prove the lemma. ∎
Now we are ready to prove Theorem 4.9.
Theorem F.6 (Theorem 4.9 in the main paper).
For the problem class $\mathcal{F}_{2}(L,\sigma,\tau)$, with the following conditions on the step sizes:
$\begin{split}\eta KL&\leq\mathcal{O}\left(\frac{1}{\beta}\wedge\frac{1}{\sqrt{MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\wedge\frac{1}{\beta\gamma LT}\right),\\ \gamma&\leq\mathcal{O}\left(\frac{\beta}{L}\right),\end{split}$ (131)
the iterates of Local SGD-M satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{\tau\sigma^{2}}{MK}\right.\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (132)
Proof.
By rearranging the terms in Lemma F.1:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}.\end{split}$ (133)
Using the condition on $\gamma$ and replacing $\frac{1}{T}\sum_{t=0}^{T-1}\Xi_{t}$ by Lemma F.5 proves the theorem. ∎
From Theorem F.6, we can derive directly the communication and sample complexity of Local SGD-M.
Corollary F.7 (Corollary 4.10 in the main paper).
Under the conditions of Theorem F.6, the required number of local steps and communication rounds for Local SGD-M to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are:
  1. (i)
    If the system-level Markov chain $X$ is stationary ($Q=\pi$),
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right).$ (134)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary ($Q\neq\pi$), with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (135)
Proof.
The proof is a direct consequence of Theorem F.6 and follows exactly the same steps in the proof of Corollary D.3. ∎
Appendix G Additional experiments
The following experiments investigate additional aspects of the scenario described in Sec. 5.
Impact of Heterogeneity: We complete the experiments in Sec. 5. In Figures 34, we plot the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for Minibatch SGD, Local SGD, and Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, the number of clients $M=\in\left\{10,100\right\}$, the number of local steps $K\in\{10,100,1000\}$, and the mixing time for every client $\tau_{m}\in\left\{10,1000\right\}$. As predicted by our analysis, the performance Local SGD stalls when the number of local updates increases, while Minibatch SGD and Local SGD-M are more robust to the heterogeneity. In particular, performance is improved as the number of samples $K$ in each communication round is increased.
Impact of Number of Clients $M$: We complete the experiments in Sec. 5. Figure 5 plots the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ for Minibatch SGD, Local SGD, and Local SGD-M, and varying numbers of clients. The algorithm parameters are given by: $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, and the number of local steps $K=100$. Observe that the gradient norm for all three algorithms decreases when the number of clients increases, indicating that collaboration is beneficial even with heterogeneous Markovian data.
Effect of mixing time: In Figure 6, we plot the gradient norm of the last iterate of Minibatch SGD, Local SGD, and Local SGD-M, with $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $\lambda=0.01$, $M\in\left\{10,100\right\}$, and the number of local steps $K\in\left\{10,100,1000\right\}$. As predicted by our analysis, the performance of Minibatch SGD, Local SGD, and Local SGD-M degrades when the mixing time increases.
Dependent vs Independent Batches: The Markovian data streams detailed in Sec. 3.2 and investigated numerically in Sec. 5 focused on the scenario where the data stream $X_{m}$ for each client $m$ consisted of samples drawn from a time-homogeneous transition kernel $P_{m}$ and initial distribution $Q$. This scenario models situations where communication is performed every $K$ collected samples. On the other hand, samples may arrive during communication, which cannot be used in client updates. For sufficiently long communication periods, the consecutive batches of $K$ samples used for updates will be approximately independent. In the stationary case, the initial distribution $\mathcal{Q}=\pi$, and it can be shown that the analysis in Sec. 4 also applies to this scenario. We therefore carry out additional experiments to compare the performance of independent batches with batches drawn sequentially from the time-homogeneous data stream described in Sec. 3.2. We highlight that the data streams of different clients remain independent.
In Figure 7, we plot the gradient norm trajectory $\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}$ of Minibatch SGD, Local SGD, and Local SGD-M, with algorithm parameters $\gamma=0.01,\eta=0.001$, $\beta=0.1$, $100$ clients, the number of local steps $K\in\{10,100,1000\}$, and the mixing time for every client $\tau_{m}=100$. We consider the following scenarios:
  • Dependent Batches: All $TK$ samples observed by client $m$ are drawn sequentially from the Markov chain $X_{m}$ with transition kernel $P_{m}$ and stationary initial distribution $\mathcal{Q}=\pi$.
  • Independent Batches: At the beginning of each communication round, the samples $x_{t}^{m}$ are drawn from the Markov chain with transition kernel $P_{m}$, and initial state $x_{t}^{m,1}\sim\pi$ draw independently of $x_{t-1}^{m}$. As such, $x_{t}^{m}$ is independent of $x_{t-1}^{m}$, while the samples $(x_{t}^{m,k})_{k\in[K]}$ remain statistically dependent.
The experimental results in Figure 7 demonstrate that Minibatch SGD, Local SGD, and Local SGD-M, perform similarly for both dependent and independent batch sampling. This supports the validity of our analysis in both the dependent and independent batch scenarios.
(a) $M=10,K=10$
(b) $M=10,K=100$
(c) $M=10,K=1000$
(d) $M=100,K=10$
(e) $M=100,K=100$
(f) $M=100,K=1000$
Figure 3: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M with $\tau_{m}=10$ for all $m\in[M]$, and for different number of clients and number of local steps.
(a) $M=10,K=10$
(b) $M=10,K=100$
(c) $M=10,K=1000$
(d) $M=100,K=10$
(e) $M=100,K=100$
(f) $M=100,K=1000$
Figure 4: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD & Local SGD-M with $\tau_{m}=1000$ for all $m\in[M]$, and for different number of clients and number of local steps.
(a) Local SGD
(b) Minibatch SGD
(c) Local SGD-M
Figure 5: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD and Local SGD-M, for $K=100$ and different numbers of clients.
(a) $M=10,K=10$
(b) $M=10,K=100$
(c) $M=10,K=1000$
(d) $M=100,K=10$
(e) $M=100,K=100$
(f) $M=100,K=1000$
Figure 6: The gradient norm of the last iterate of 3 algorithms as a function of the mixing time with different numbers of clients and different numbers of local steps.
(a) Local SGD
(b) Minibatch SGD
(c) Local SGD-M
(d) Local SGD
(e) Minibatch SGD
(f) Local SGD-M
(g) Local SGD
(h) Minibatch SGD
(i) Local SGD-M
Figure 7: Gradient norm as a function of the number of communication rounds for Local SGD, Minibatch SGD and Local SGD-M with $100$ clients, mixing times $\tau_{m}=100$ for all $m\in[M]$, number of local steps $K=10$ (first row), $K=100$ (second row) and $K=1000$ (third row).